OpenStudy (anonymous):

Write the expression as a single trigonometric function: cos15cos75+sin15sin75

OpenStudy (chaise):

cos60 or sin30

OpenStudy (unklerhaukus):

\[\cos (u)\cos (v)\mp\sin (u)\sin (v)=\cos (u\pm v)\]

OpenStudy (anonymous):

I'm not sure answer , do u missing i sin?

OpenStudy (anonymous):

\[\text{Cos}[15{}^{\circ}]\text{Cos}[75{}^{\circ}]+\text{Sin}[15{}^{\circ}]\text{Sin}[75{}^{\circ}]=\frac{1}{2} \]